OpenStudy (anonymous):

Solve the system, using substitution. Write the solution as an ordered pair (m, n). 2m - 3n = 6 and m = 3 - 2n A. (6, 2) B. (3, 0) C. (-3, -4) D. (0, 2) E. no solution I put C is this correct?

OpenStudy (anonymous):

i got B for this one

OpenStudy (watchmath):

Substitute the second to the first \(2(3-2n)-3n=6\) \(6-4n-3n=6\) \(-7n=0\) \(n=0\) Hence \(m=3-0=3\) So (3,0) is the answer

OpenStudy (anonymous):

Hmm alright do you guys all agree?

OpenStudy (anonymous):

Yep well u do.. okay (: Thanks!!